# Gueter Josmy Faure1    Min-Hung Chen2    Jia-Fong Yeh1    Hung-Ting Su1 Shang-Hong Lai4    Winston H. Hsu1,3

### v3

@@ Abstract
Existing research often treats long-form videos as extended short videos, leading to several limitations: inadequate capture of long-range dependencies, inefficient processing of redundant information, and failure to extract high-level semantic concepts. To address these issues, we propose a novel approach that more accurately reflects human cognition. [^1] We elaborate on this in Section A.7 This paper introduces HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics, a model that simulates episodic memory accumulation to capture action sequences and reinforces them with semantic knowledge dispersed throughout the video. Our work makes two key contributions: First, we develop an Episodic COmpressor (ECO) module that efficiently aggregates crucial representations from micro to semi-macro levels, overcoming the challenge of long-range dependencies. Second, we propose a Semantics reTRiever (SeTR) that enhances these aggregated representations with semantic information by focusing on the broader context, dramatically reducing feature dimensionality while preserving relevant macro-level information. This addresses the issues of redundancy and lack of high-level concept extraction. Extensive experiments demonstrate that HERMES achieves state-of-the-art performance across multiple long-video understanding benchmarks in both zero-shot and fully-supervised settings. Our code will be made public.
Figure 1: Semantic Knowledge and Episodic Memory Aggregation: Our Episodic COmpressor (ECO) processes input video frames and aggregates key episodes: (I) A cosmic setting with a planet and starfield, (II) A narrator explaining the scene, and (III) An aircraft viewed from inside and out. Concurrently, our Semantics reTRiever (SeTR) identifies high-level semantic cues throughout the video, including: (I) The theme of space exploration and (II) Human interaction with technology. This dual-level approach enables comprehensive video understanding by capturing both specific events and overarching concepts.
@@ 1 Introduction
Long-form video understanding is challenging for several reasons. First and foremost is the temporal complexity, as it requires handling a large number of frames throughout the video. Second, it requires a semantic understanding of high-level concepts as well as the narrative structure. The third challenge is the memory and computational constraints, making it non-trivial to solve the previous two challenges. Attempts to address these issues have been made by researchers who mainly borrow ideas from short videos (Wu & Krahenbuhl, 2021; Miech et al., 2020), which is a more mature area of research encompassing action recognition and video classification, among others, and for which datasets are more abundant. However, these approaches, which adopt techniques such as pooling (Faure et al., 2023), or 3D convolutions (Feichtenhofer et al., 2019), often do not fully account for the unique characteristics of long videos that distinguish them from a simple concatenation of short video segments. Some ideas about short-video modeling, especially for those at the spatial level, may also hold for longer ones, but when it comes to long-term modeling, macro-level representations should be extracted efficiently.
In video understanding, we can distinguish between two types of information: episodic and semantic. Episodic information refers to specific, sequential events that occur in a video, while semantic information encompasses overarching themes and concepts. To illustrate, imagine walking through a scene at a birthday party. Episodic information might include observing five people singing a birthday song, followed by one person cutting a cake. These are specific, time-bound events. In contrast, semantic information might involve recognizing decorations scattered throughout the scene, instantly comprehending that you’re witnessing a birthday party. This high-level understanding provides a concise overview of the scene and actions, transcending specific moments. Building on these concepts, we propose temporal-coHERent long-forM understanding with Episodes and Semantics (HERMES), a model designed to capture both episodic and semantic information in long-form videos. HERMES  comprises two key components: ECO  and SeTR. The Episodic COmpressor (ECO) aggregates key contiguous information as we process the video, shaping the model’s understanding of the scene sequentially without cluttering it. Complementing this, the SEmantic reTRiever (SeTR) identifies and extracts high-level cues that provide a concise overview of the scene and actions. HERMES  achieves state-of-the-art performance on four long-form video understanding benchmarks in both zero-shot and fully-supervised settings, notably outperforming the state-of-the-art by a significant 7.3% on LVU(Wu & Krahenbuhl, 2021) and 14.9% on MovieChat-1k (Song et al., 2024).
@@
  • We address the challenge of efficiently processing and understanding long-form videos, a problem that has been hindered by computational constraints and the difficulty of capturing complex temporal dynamics over extended durations.
  • We propose an Episodic COmpressor (ECO) to stream through the video and keep important episodes by aggregating similar scenes. ECO enables efficient processing of extended video sequences while preserving temporal coherence and narrative structure.
  • We develop a Semantics reTRiever (SeTR) that enhances the model’s understanding of long videos by distilling high-level semantic cues, providing a cohesive framework for understanding the context and themes within long-form videos.
Through comprehensive evaluation across multiple benchmarks and detailed ablation studies, we validate the effectiveness of ECO and SeTR, and demonstrate their complementary roles in enhancing long-form video understanding
2 Problem Setting
Given a long video $V=\{f_{1},f_{2},\ldots,f_{N}\}$, where $f_{i}$ represents the $i$-th frame and $N$ is the total number of frames, our objective is to develop a model $\mathbf{M}$ that can efficiently process $V$ and construct an internal understanding $U$ of its content. This understanding should enable the model to answer queries $Q$ or follow instructions $I$ related to the video content. Formally, we aim to optimize the function:
@@ 2 Problem Setting
  • $U$ captures both episodic and semantic information from $V$.
  • $U$ can be used to maximize the probability $P(A|Q,U)$ of generating correct answers $A$ to queries $Q$ about the video content.
@@
    Temporal Complexity: Efficiently processing $N$ frames, where $N$ can be very large (thousands of frames for minutes-long videos).
@@
    Semantic Understanding: Extracting high-level concepts and narrative structure from the video content.
@@
    Memory Constraints: Developing a method that can maintain relevant information without exhausting computational resources.
We aim to address these challenges, by proposing two key contributions:
@@
    Episodic COmpressor (ECO):
    $ECO:{f_{1},f_{2},\ldots,f_{N}}\rightarrow{e_{1},e_{2},\ldots,e_{K}}$ (2)
    where $K\ll N$, and $e_{i}$ represents a compressed episodic memory.
@@
    Semantics reTRiever (SeTR):
@@
    where $L\ll N$, and $s_{i}$ represents extracted semantic knowledge.
The final understanding $U$ is then generated by combining the outputs of ECO and SeTR:
@@
where $G$ is a function that integrates episodic and semantic information. This formulation allows us to approach long-form video understanding in a way that more closely mimics human cognition, addressing the identified challenges while maintaining computational efficiency.
3 Proposed Framework: HERMES
This paper is not about a new LLM or a new way to fine-tune existing LLMs or VLMs. It focuses on leveraging what we know about how humans understand visual scenes to guide the model through the same process. Although this work uses an LLM for autoregressive prediction, the core ideas of episodic memory compression (ECO) and semantic knowledge retrieval (SeTR) can be applied to other models where learning contiguous sequences and high-level representations is advantageous.
@@ 3.1 Window Encoder
1:$\mathcal{A}\leftarrow\mathcal{M}\oplus F_{w}$
@@
3.2 ECO: Episodic COmpressor
The proposed Episodic COmpressor (ECO)  aggregates the video frames into episodes. This module maintains a memory buffer with a maximum number of episodes $E$. Upon receiving a window of $F_{w}$ frame features, we first check whether the buffer $\mathcal{M}$ has sufficient bandwidth to support the incoming features. If it does, we simply concatenate them to the buffer; otherwise, we proceed with the compression. At its core, ECO is a distribution process that determines the episode to which a certain frame belongs. It can be summarized as:
=$\mathcal{M}=\begin{cases}\mathcal{M}\oplus F_{w}&\text{if }\|\mathcal{M}\|+\|F_{w}\|\leq E\\ \text{ECO}(\mathcal{M},F_{w})&\text{otherwise}\end{cases}$ (5)
Where $\oplus$ is the concatenation operation, $\|\mathcal{M}\|$ and $\|F_{w}\|$ the sizes of the buffer and the incoming features, respectively.
ECO works as Algorithm 1. As long as concatenating the new window and the buffer results in a size greater than $E$, we compute the cosine similarity between each pair of frame features in $\mathcal{M}\oplus F_{w}$. We then iteratively merge the most similar frames until the size constraint $E$ is satisfied. Specifically,
where $\mathcal{M}$ is the existing buffer, $F_{w}$ represents the incoming window of $w$ frame features, $\mathcal{A}$ is the concatenated buffer and new window, and $\|\mathcal{A}\|$ is the size of $\mathcal{A}$. To summarize Algorithm 1, $\frac{\mathcal{A}_{i}\cdot\mathcal{A}_{j}}{\|\mathcal{A}_{i}\|\|\mathcal{A}_{j}\|}$ computes the cosine similarity between frame features $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$, $\arg\max_{i\neq j}$ finds the pair of frames with the highest cosine similarity, $\frac{(\mathcal{A}_{i^{*}}+\mathcal{A}_{j^{*}})}{2}$ combines the most similar frames, and $\mathcal{A}\setminus\mathcal{A}_{j^{*}}$ removes the frame $\mathcal{A}_{j^{*}}$ from $\mathcal{A}$ after merging. The process repeats until the size of $\mathcal{A}$ is within the maximum allowed episodes $E$.
Similarities can be drawn with He et al. (2024), where cosine similarity serves as the basis for frame reduction. However, their approach is notably inefficient and less intuitive. For a buffer of size $S$, they iterate $S$ times until the buffer reaches capacity, after which each new incoming frame is compared against every other frame in the buffer.
@@ 3.3 Episodic Q-Former
The Episodic Q-Former uses the same architecture as the original Q-Former (Li et al., 2023a) and is loaded with weights pretrained by Dai et al. (2023). However, we insert ECO as a pruning module within the Q-Former to combine and branch queries into episodes over the long video. Given initial queries and instructions, we perform self-attention on these queries and then cross-attention between the queries and the visual representations $\mathcal{M}$. The enhanced queries then undergo an ECO-like process, where we iteratively merge similar queries across video windows, effectively forming video query episodes of high information density. The following equation summarizes the process,
@@ 3.4 SeTR: Semantics reTRiever
This process effectively reduces the number of frames from $N$ to $\frac{N}{k}$, consolidating semantic information while maintaining the most relevant features across the video timeline. The resulting semantic representations are denoted as $F^{\prime}\in\mathbb{R}^{B\times\frac{N}{k}\times T\times C}$. We evaluate the effectiveness of this approach in Section 4.3. While ToMe (Bolya et al., 2022) have explored token reduction in vision transformers, their approach and objectives differ significantly from ours. Their method focuses on minor token reductions within individual frames, specifically between different layers of a Vision Transformer. In contrast, SeTR retains the most salient frames while significantly reducing redundancies.
@@ 3.6 From Representations to Natural Language
After obtaining the episodic representations $Q$ and the semantic representations $Q_{sem}$, we prepare them for input into a Large Language Model (LLM). Specifically, we concatenate $Q$ and $Q_{sem}$ to form a unified representation vector. This concatenated vector is then projected into the input embedding space of the LLM using a learned linear transformation. In our implementation, we utilize a Vicuna-7B model (Chiang et al., 2023) as LLM. The model, conditioned on this projected representation and guided by task-specific instructions, generates the requested natural language output. This approach allows us to leverage the LLM’s pretrained knowledge and language generation capabilities while incorporating our task-specific episodic and semantic information. The process is summarized by the following equation:
$\hat{Y}=\text{LLM}(W[Q;Q_{sem}]+b,I)$ (8)
where $\hat{Y}$ is the generated output, $[Q;Q_{sem}]$ denotes the concatenation of $Q$ and $Q_{sem}$, $W$ and $b$ are the learned projection matrix and bias respectively, and $I$ represents the task-specific instructions.

### v4

# HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics

@@ Abstract
Long-form video understanding presents unique challenges that extend beyond traditional short-video analysis approaches, particularly in capturing long-range dependencies, processing redundant information efficiently, and extracting high-level semantic concepts. To address these challenges, we propose a novel approach that more accurately reflects human cognition. This paper introduces HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics, featuring two versatile modules that can enhance existing video-language models or operate as a standalone system. Our Episodic COmpressor (ECO) efficiently aggregates representations from micro to semi-macro levels, reducing computational overhead while preserving temporal dependencies. Our Semantics reTRiever (SeTR) enriches these representations with semantic information by focusing on broader context, dramatically reducing feature dimensionality while preserving relevant macro-level information. We demonstrate that these modules can be seamlessly integrated into existing SOTA models, consistently improving their performance while reducing inference latency by up to 43% and memory usage by 46%. As a standalone system, HERMES achieves state-of-the-art performance across multiple long-video understanding benchmarks in both zero-shot and fully-supervised settings. Our project page and code can be found here.
Figure 1: Semantic Knowledge and Episodic Memory Aggregation: Our Episodic COmpressor (ECO) processes and aggregates temporal information across different scales: (I) Social interactions among adolescents in an outdoor setting, and (II) Complex family dynamics portrayed through parent-child interactions. Simultaneously, our Semantics reTRiever (SeTR) extracts high-level semantic information: (I) The contextual environment of a baseball game, and (II) The intersection of media consumption and domestic life through a news broadcast. This dual-level approach enables comprehensive video understanding by capturing both specific events and overarching concepts.
@@ 1 Introduction
Long-form video understanding is challenging for several reasons. First and foremost is the temporal complexity, as it requires handling a large number of frames throughout the video. Second, it requires a semantic understanding of high-level concepts as well as the narrative structure. The third challenge is the memory and computational constraints, making it non-trivial to solve the previous two challenges. Attempts to address these issues have been made by researchers who mainly borrow ideas from short videos [43, 24], which is a more mature area of research encompassing action recognition and video classification, among others, and for which datasets are more abundant. However, these approaches, which adopt techniques such as pooling [11], or 3D convolutions [12], often do not fully account for the unique characteristics of long videos that distinguish them from a simple concatenation of short video segments. Some ideas about short-video modeling, especially for those at the spatial level, may also hold for longer ones, but when it comes to long-term modeling, macro-level representations should be extracted efficiently.
In video understanding, we can distinguish between two types of information: episodic and semantic [^1] We elaborate on this in Section H.9. Episodic information refers to specific, sequential events that occur in a video, while semantic information encompasses overarching themes and concepts. To illustrate, consider the scene presented in Figure 1. Episodic information includes observing adolescents interacting at a baseball game, followed by tense exchanges between a mother and father. These are specific, time-bound events that unfold sequentially. In contrast, semantic information involves recognizing the broader context of youth sports culture and the backdrop of media’s influence on domestic life. This high-level understanding concisely overviews the scene and actions, transcending specific moments.
Building on these concepts, we propose temporal-coHERent long-forM understanding with Episodes and Semantics (HERMES), featuring two modular components that can either work together as a complete system or integrate into existing models. The Episodic COmpressor (ECO) aggregates key contiguous information as we process the video, shaping understanding sequentially while reducing computational overhead and the SEmantic reTRiever (SeTR) identifies and extracts high-level cues that provide a concise overview of the scene and actions. HERMES  achieves state-of-the-art performance on four long-form video understanding benchmarks in both zero-shot and fully-supervised settings, outperforming the state-of-the-art by 7.3% on LVU[43] and 14.9% on MovieChat-1k [32].
@@
  • We develop a versatile framework for processing and understanding long-form videos that can either operate as a standalone system or enhance existing models through modular integration.
  • We propose an Episodic COmpressor (ECO) that can replace or augment existing memory mechanisms, consistently improving model performance while reducing inference latency and GPU memory usage by up to 43% and 46%, respectively.
  • We develop a Semantics reTRiever (SeTR) that enhances video understanding by distilling high-level semantic cues, providing substantial accuracy improvements with minimal computational overhead.
Through comprehensive evaluation across multiple benchmarks, detailed ablation studies, and extensive integration experiments with existing SOTA models, we validate the effectiveness of ECO and SeTR, demonstrating their complementary roles in enhancing long-form video understanding both as standalone components and as plug-in modules.
2 Related Work
Action recognition is an essential task in video understanding, primarily focusing on identifying specific actions within short video clips. Various approaches have been developed, with convolutional neural networks forming the core of many of them. Early work by [17] utilized 3D convolutions, while [40] employed temporal convolutions. 2D CNNs coupled with temporal modeling have also been explored, with representative works such as Temporal Difference Networks (TDN) [25] and Event Adaptive Networks (EAN) [35]. More recently, transformer-based models have gained prominence with works such as [11], [45], and [48].
Video question answering (VideoQA) aims to answer questions related to video content, requiring a deep understanding of both visual and textual information. Datasets such as ActivityNet-QA [46] for short videos, and MovieChat-1k for long videos [32] provide benchmarks for evaluating models in this field, allowing for several research endeavors on this subject [50, 52, 27].
Long-form video understanding presents unique challenges due to the extended duration and complex narrative structures involved. Datasets with these properties include LVU [43], COIN [34], Breakfast [18], and MovieChat-1k [32]. Traditional approaches to tackling such a task often extend methods designed for short videos to handle longer sequences, such as pooling over the temporal dimension [33, 11]. Other methods such as [43, 14, 44] and [32] explore memory techniques via token compression. Additionally, [36] introduced a video semantic compression framework using low-level bitrate coding. [41] introduced selective structured state-spaces for long-form videos, followed by others [15, 16] exploiting the ability of state-space models to retain long-term context.
Video-Language Models: Recent advancements in large language models (LLMs) [37, 6] have piqued researchers’ curiosity regarding their use for video understanding [22]. It turns out to be a good match, as understanding videos often involves transforming their content into words, whether it’s video captioning, video question answering, or even action classification. Frameworks such as [32] and [14] employ memory techniques to handle extensive video content while [29] presents TimeChat, explicitly conditioning the model to manage time-dependent information.
3 Problem Statement
Given a long video $V=\{f_{1},f_{2},\ldots,f_{N}\}$, where $f_{i}$ represents the $i$-th frame and $N$ is the total number of frames, our objective is to develop a model $\mathbf{M}$ that can efficiently process $V$ and construct an internal understanding $U$ of its content. This understanding should enable the model to answer queries $Q$ or follow instructions $I$ related to the video content. Formally, we aim to find an optimal function:
@@ 3 Problem Statement
  • $U$ captures episodic and semantic information from $V$.
  • $U$ can be used to maximize the probability $P(A|Q,U)$ of generating correct answers $A$ to queries $Q$ about the video.
@@
    Temporal Complexity: Efficiently processing $N$ frames, where $N$ can be very large.
@@
    Semantic Understanding: Extracting high-level concepts and narrative structure from video content.
@@
    Memory Constraints: Developing a method to maintain relevant information without exhausting computational resources.
Addressing these challenges requires an approach that can effectively compress temporal information while preserving both detailed episodic content and high-level semantic understanding. In the following section, we propose a cognitively inspired framework to tackle these challenges.
4 Proposed Framework: HERMES
Our goal is to enhance video understanding by loosely drawing inspiration from human visual processing, rather than developing a new LLM or fine-tuning existing ones. To achieve this, we introduce a method that, given a video and a set of instructions, generates the specified output, such as video question answering (VQA) or video classification. Figure 2 provides a high-level overview of our framework.
Our approach addresses the challenges identified in Section 3 through two core principles of human perception:
@@ 4 Proposed Framework: HERMES
    An Episodic COmpressor (ECO), which structures a video into meaningful segments:
    $ECO:\{f_{1},f_{2},\ldots,f_{N}\}\rightarrow\{e_{1},e_{2},\ldots,e_{K}\}$ (2)
    where $K\ll N$, and $e_{i}$ represents compressed episodes.
@@
    A Semantics reTRiever (SeTR), which extracts high-level semantic context:
@@
    where $L\ll N$, and $s_{i}$ represents extracted semantics.
The final understanding $U$ is generated by combining the outputs of ECO and SeTR:
@@
where $G$ is a function that integrates episodic and semantic information.
@@ 4.1 Window Encoder
4.2 ECO: Episodic COmpressor
Long videos often contain redundant information, making it crucial to identify and consolidate key episodic elements efficiently. To address this, we propose ECO which maintains a memory buffer with a maximum number of episodes $E$. Upon receiving a window of frame features, $\mathcal{W}_{k}$, we first check whether the buffer $\mathcal{M}$ has sufficient bandwidth to support the incoming features. If it does, we simply concatenate them to the buffer; otherwise, we proceed with the compression. At its core, ECO is a distribution process that determines the episode to which a certain frame belongs. It can be summarized as:
=$\mathcal{M}=\begin{cases}\mathcal{M}\oplus\mathcal{W}_{k}&\text{if }\|\mathcal{M}\|+\|\mathcal{W}_{k}\|\leq E\\ \text{ECO}(\mathcal{M},\mathcal{W}_{k})&\text{otherwise}\end{cases}$ (5)
where $\oplus$ is the concatenation operation, $\|\mathcal{M}\|$ and $\|\mathcal{W}_{k}\|$ are the sizes of the buffer and the incoming features, respectively.
@@ 4.2 ECO: Episodic COmpressor
1:$\mathcal{A}\leftarrow\mathcal{M}\oplus\mathcal{W}_{k}$
@@
ECO works as Algorithm 1 where $\mathcal{M}$ is the existing buffer, $\mathcal{W}_{k}$ represents the incoming window of frame features, $\mathcal{A}$ is the concatenated buffer and new window, and $\|\mathcal{A}\|$ is the size of $\mathcal{A}$. To summarize ECO, $\frac{\mathcal{A}_{i}\cdot\mathcal{A}_{j}}{\|\mathcal{A}_{i}\|\|\mathcal{A}_{j}\|}$ computes the cosine similarity between frame features $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$, $\arg\max_{i\neq j}$ finds the pair of frames with the highest cosine similarity, $\frac{(\mathcal{A}_{i^{*}}+\mathcal{A}_{j^{*}})}{2}$ combines the most similar frames, and $\mathcal{A}\setminus\mathcal{A}_{j^{*}}$ removes the frame $\mathcal{A}_{j^{*}}$ from $\mathcal{A}$ after merging. The process repeats until the size of $\mathcal{A}$ is within the maximally allowed episodes $E$.
@@ 4.3 Episodic Q-Former
To aggregate learned queries into episodes as we did the video features, we integrate ECO as a pruning module within the Q-Former architecture (initialized with weights from [7]). Given initial queries and instructions, we perform self-attention on these queries followed by cross-attention between the queries and visual representations $\mathcal{M}$. The enhanced queries then undergo an ECO-like process, where we iteratively merge similar queries across video windows, effectively forming video query episodes of high information density. The following equation summarizes the process,
@@ 4.4 SeTR: Semantics reTRiever
This process effectively reduces the number of frames from $N$ to $\frac{N}{k}$, consolidating semantic information while maintaining the most relevant features across the video timeline. The resulting semantic representations are denoted as $F^{\prime}\in\mathbb{R}^{B\times\frac{N}{k}\times T\times C}$. We evaluate the effectiveness of this approach in Section 5.4. While ToMe [3] have explored token reduction in vision transformers, their approach and objectives differ significantly from ours. Their method focuses on minor token reductions within individual frames, specifically between different layers of a Vision Transformer. In contrast, SeTR retains the most salient frames while significantly reducing redundancies.
@@ 4.6 From Representations to Natural Language
After obtaining the episodic representations $Q$ and the semantic representations $Q_{sem}$, we prepare them for input into a Large Language Model (LLM). Specifically, we concatenate $Q$ and $Q_{sem}$ to form a unified representation vector. This concatenated vector is then projected into the input embedding space of the LLM using a learned linear transformation. In our implementation, we utilize a Vicuna-7B model [6] as the LLM. The model, conditioned on this projected representation and guided by task-specific instructions, generates the requested natural language output. This approach allows us to leverage the LLM’s pretrained knowledge and language generation capabilities while incorporating our task-specific episodic and semantic information. The process is summarized by the following equation:
$\hat{Y}=\text{LLM}(U,I)$ (8)
where $U=W[Q;Q_{sem}]+b$ denotes the understanding stemming from the aggregation of semantic and episodic information, $\hat{Y}$ is the generated output, $[Q;Q_{sem}]$ the concatenation of $Q$ and $Q_{sem}$, $W$ and $b$ are the learned projection matrix and bias respectively, and $I$ represents the task-specific instructions.
Model Global Breakpoint
Acc. Score Acc. Score
MovieChat [32] 63.7 3.15 48.1 2.46
Video-ChatGPT [22] 58.7 2.89 47.8 2.43
Video-LLaMA [49] 56.3 2.72 45.8 2.11
VideoChat [20] 60.2 3.08 46.3 2.32
HERMES (Ours) 78.6 4.23 57.3 3.29
HERMES (Ours) 84.9 4.40 65.8 3.65
LVU
Model Content Metadata Avg Breakfast COIN
Relation Speak Scene Director Genre Writer Year
FACT [21] - - - - - - - - 86.1 -
Obj. Transformer [43] 53.1 39.4 56.9 52.1 54.6 34.5 39.1 47.1 - -
VIS4mer [15] 57.1 40.8 67.4 62.6 54.7 48.8 44.8 53.7 88.2 88.4
TranS4mer [16] 59.5 39.2 70.9 63.9 55.9 46.9 45.5 54.5 90.3 89.2
S5 [41] 67.1 42.1 73.5 67.3 65.4 51.3 48.0 59.2 90.7 90.8
Movies2Scenes [5] 71.2 42.2 68.2 70.9 57.8 55.9 53.7 60.0 - -
MA-LMM [14] 58.2 44.8 80.3 74.6 61.0 70.4 51.9 63.0 93.0 93.2
HERMES (Ours) 67.6 47.5 90.0 82.6 69.5 77.2 57.7 70.3 95.2 93.5
Table 2: SOTA Comparison on the LVU, Breakfast and COIN datasets: The table presents Top-1 accuracy for various models. Unlike the minor incremental improvements observed among other methods, our model demonstrates a significant performance leap, outperforming its nearest competitor by 7.3% on LVU and 2.2% on Breakfast. The highest score is highlighted in bold, and the second highest is underlined.
Model Acc. Time Mem. (GB)
LongVA (7B) 54.11 1 42.5
+ ECO 54.19 0.700 (-30%) 22.9
+ SeTR 54.56 0.726 (-27%) 32.7
Table 3: Zero-shot performance comparison of LongVA with and without ECO and SeTR integration on VideoMME.
Model Acc. Time Mem. (GB)
LLaVA-OV (7B) 58.26 1 40.6
+ ECO 58.93 0.650 (-35%) 33.4
+ SeTR 59.30 0.673 (-33%) 33.4
Table 4: Zero-shot performance comparison of LLaVA-OneVision with and without ECO and SeTR integration on VideoMME.
Model Acc. Score Time Mem. (GB)
MA-LMM 73.3 4.05 1 30.2
+ ECO 76.7 4.14 0.569 (-43%) 30.2
+ SeTR 77.1 4.16 1.015 (+1.5%) 32.5
Table 5: Zero-shot comparison of MA-LMM with and without ECO as memory manager and integrating SeTR on MovieChat-1k.